freddy the cat is stuck in a tree, his owner has a ladder that that is 20 feet long. If the owner places the ladder 3 feet away from the tree, how high up is freddy?
Answer by richwmiller(17219)

not enough info.
We don't know if the ladder is too short, just reaches or is too long.
Assuming that it just reaches,20 is the hypotenuse and 3 is one side.
3^2+a^2=20^2
9+a^2=400
a^2=391
a=sqrt(391)
a=19.77